# Super-Resolved Canopy Height Mapping from Sentinel-2 Time Series Using LiDAR HD Reference Data across Metropolitan France

### v1

@@ 3.1 Model
In this work, we propose an end-to-end model, named THREASURE-Net, that performs super-resolution and regression tasks simultaneously. Given a Sentinel-2 SITS $s=[s_{t_{1}},s_{t_{2}},\dots,s_{T}]\in\mathbb{R}^{T\times C\times H\times W}$ acquired at dates $t=[t_{1},t_{2},\dots,t_{T}]\in\mathbb{R}^{T}$, the model aims to predict the tree height $95^{\text{th}}$ percentile corresponding to the reference date $t_{L}$, producing an output of shape $f\cdot H\times f\cdot W$, where $f$ denotes the SR factor. Note that for the sake of simplicity, we omit batch dimension in the model description.
@@
The RDB-FE module is extracted from Residual Residual Dense network (RDN) (Zhang et al., 2018), which is mainly used for natural image SISR task and originally consists of four parts: shallow feature extraction net (SFENet), residual dense blocks (RDBs), dense feature fusion (DFF) and finally the up-sampling net (UPNet). Given the multi-temporal nature of our model, at this stage, we only deploy the feature extraction modules (SFENet, RDBs, and DFF). The up-sampling module is applied later to the temporally encoded features, rather than to each image individually.
@@
Temporal feature extraction block. To extract a unique set of temporal features for each Sentinel-2 series, we exploit the Lightweight Temporal Self-Attention mechanism proposed by Garnot and Landrieu (2020). The module captures dependencies across the multi-temporal Sentinel-2 time series and allows the network to dynamically weight the contribution of each timestamp, by using keys $K$, queries $Q$ and values $V$ to produce a self-attention mask. While $K$ and $V$ are directly derived from encoded features $f$ and encoded temporal positions $p$ of Sentinel-2, $Q$ is a learnable parameter. For detailed explanation of Lightweight Temporal Transformer, the reader is invited to consult the original article. The simplified version of this algorithm can be summarized as follows (here we omit the explanation of the multi-head strategy):
Let $\mathbf{f}=[\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{T}]\in\mathbb{R}^{T\times F_{1}}$ be the sequence of encoded S2 observations acquired at times $\{t_{1},\dots,t_{i},\dots,t_{T}\}$. Each timestamp is associated with a temporal encoding $\mathbf{p}_{i}\in\mathbb{R}^{d}$ obtained using a sinusoidal positional function:
@@
where $\tau$ is a scaling constant and $k$ indexes the embedding dimensions.
$\tilde{\mathbf{f}}_{i}=\mathbf{f}_{i}+\mathbf{p}_{t_{i}}.$ (2)
The attention mechanism computes keys and values as linear projections:
$\mathbf{K}=\tilde{\mathbf{F}}W_{K},\quad\mathbf{V}=\tilde{\mathbf{F}}W_{V},$ (3)
with learnable weights $W_{K},W_{V}\in\mathbb{R}^{d\times d_{a}}$.
A single learnable query vector $\mathbf{q}\in\mathbb{R}^{d_{a}}$ is used to summarize the temporal sequence by computing attention weights:
The attention weights are then obtained as:
$\boldsymbol{\alpha}=\operatorname{softmax}\!\left(\frac{\mathbf{K}\,\mathbf{q}_{L}^{\top}}{\sqrt{d_{a}}}\right),$
giving the final aggregated temporal representation is:
$\mathbf{f}_{temp}=\boldsymbol{\alpha}^{\top}\mathbf{V}.$ (4)
with final size $T\times F_{2}\times H\times W$}, where the new feature dimension $F_{2}$ is a hyperparameter.
@@
where learnable weights $W_{Q}\in\mathbb{R}^{d\times d_{a}}$.
@@
As shown in Equations 6 and 7, the resulting DOYs represent relative time offsets (in days) from the beginning and the middle of the year, respectively. This preprocessing highlights the seasonal relationship between the LiDAR reference date and the Sentinel-2 time series, allowing the model to better capture phenological phase alignment. During inference, predictions are by default conditioned on July 1st, whereas during training the DOY representation preserves temporal offsets around this date, enabling the model to encode whether an observation occurs earlier or later in the seasonal cycle.
@@
Tree height regression. The final tree-height predictions are produced by a Multi-Layer Perceptron (MLP) that takes super-resolved features as input and outputs a regression map of tree heights at the selected resolution.
@@ 3.2 Loss functions
To optimize the model, the following loss functions are used: Mean Absolute Error (MAE) for tree height prediction, and Weighted Gradient Difference Loss (WGDL) to enhance the texture and edges of super-resolved features.
$\mathcal{L}_{total}=\omega_{height}\cdot\mathcal{L}_{height}+\omega_{wgdl}\cdot\mathcal{L}_{wgdl}$ (8)
Tree height regression Loss. As mentioned earlier, the MAE loss is employed for tree height prediction. However, we assume that most patches exhibit low variability in pixel values due to the homogeneous forest structure, which may lead to imbalance within a batch. To solve this issue, instead of directly averaging pixel-wise loss values across the batch, we first average them within each patch and then across patches in the batch as follows:
$\mathcal{L}_{\text{batch}}=\frac{1}{B}\sum_{b=1}^{B}\left(\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\ell(p_{b,i})\right)$ (9)
where $B$ is the batch size, $N_{b}$ is the number of valid pixels within a patch and $\ell(p_{b,i})$ are pixel-wise loss values.
@@
Note that the masked pixels are excluded from the loss.
@@ 4.1.1 Evaluation Metrics
To evaluate the performance of the model, we report four complementary metrics: the coefficient of determination (R²), the MAE, the root mean squared error (RMSE) and the relative mean absolute error (rMAE). R² measures how well the predicted values follow the variability of the true heights, with higher values indicating better explanatory power. MAE reflects the average magnitude of the prediction errors, while RMSE emphasizes larger errors more strongly and, therefore, highlights occasional large deviations. Finally, rMAE expresses the average error relative to the true values, allowing comparison across height ranges with different scales.
@@ 4.1.2 Model Configuration
We evaluate THREASURE-NET at three spatial resolutions: 10 m, 5 m, and 2.5 m. Only pixels labeled as vegetation are used for training – specifically, permanent vegetation taller than 1.5 m. A mask is applied to exclude non-vegetation pixels.
@@
Finally, the prediction head consists of an MLP with layer sizes (64, 128, 64) that outputs the final single-channel height estimate.
The maximum time-series length is set to 12, and all series shorter than 5 observations are discarded. During training, a random sampling strategy is applied when the time-series length exceeds 12, whereas an equal-range sampling strategy is used during validation to ensure consistency of the validation data.
The input Sentinel-2 patches contain $64\times 64$ pixels ($640\times 640$ m²), augmented with border margins that are passed through the model but discarded from the output before loss computation and inference, in order to avoid border effects due to 2D convolution operations. During training, we apply a random sampling strategy within the borders of each LiDAR HD patch ($1\times 1$ km²), whereas during validation the central region of the patch is systematically selected. For this reason, the initial Sentinel-2 patches were constructed with 250 m margins on each side relative to the LiDAR patches.
As mentioned in Section 2, MAE and WGDL losses are used for model optimization. All input features were standardized to ensure stable training.
@@
Figure 4: Prediction–target scatterplots for our 10 m, 5 m, and 2.5 m models. For visualization purposes, we randomly sampled an equal number of points ($\approx$96 million) from each model, corresponding to the total number of points in the 10 m dataset. The identity line is shown in white. Note that although the minimum target values in the dataset are set to 1.5 m, we report the model predictions without applying any clipping.
4.2 Results and Discussion
4.2.1 Quantitative Results
Table 1: Quantitative results for our models of 10, 5 and 2.5m resolutions. We also report the performance of two other recent state-of-the-art algorithms: FORMS-T (Schwartz et al., 2025) at 10m and Open-Canopy (Fogel et al., 2025) at 1.5m resolution. N/R stands for “Not Reported”.
Model Res., m MAE, m MAE, % R2 RMSE, m
THREASURE-Net 10 2.62 22.0 0.75 3.56
5 2.72 25.4 0.72 3.70
2.5 2.88 30.0 0.67 3.91
Open-Canopy 1.5 2.52 22.9 N/R 4.02
FORMS-T 10 3.06 N/R 0.68 N/R
In this subsection, we report the qualitative results of our model performance at different resolutions. The different metrics are reported in Table 1, along with the performance of two other recent state-of-the-art algorithms: FORMS-T and Open-Canopy, taken from the original articles. FORMS-T (Schwartz et al., 2025) uses single-date Sentinel-1 and Sentinel-2 imagery, together with GEDI LiDAR data as reference, to predict canopy structure at 10 m resolution. Open-Canopy (Fogel et al., 2025), in contrast, relies on a single VHR SPOT-6/7 image combined with LiDAR HD to map fine-scale canopy height at 1.5 m.
@@ 4.2.1 Quantitative Results
Although Open-Canopy obtains the lowest MAE overall, its RMSE is higher than the RMSE achieved by THREASURE-Net at each of the three resolutions. This indicates that Open Canopy performs well on average errors but struggles more with extreme values, tending to regress predictions toward the mean. In our case, quantitative performance decreases as the target resolution becomes finer. Nevertheless, the 2.5 m version of our model remains competitive with Open-Canopy, despite relying exclusively on freely available 10 m Sentinel-2 imagery. By contrast, Open-Canopy is trained on 1.5 m SPOT-6/7 data, which are commercial products or accessible only through specific research agreements. This highlights an important advantage of our approach: comparable performance can be achieved without relying on costly high-resolution remote sensing input data. Moreover, all variants of THREASURE-Net reached better results than 10 m resolution FORMS-T, despite relying on Sentinel-2 data only, whereas FORMS-T uses Sentinel-1 and Sentinel-2 data, which requires considerably heavier data processing. Figure 4 presents the prediction–target scatterplots for our models. The 10 m model produces predictions that cluster closely around the identity line, indicating strong agreement with the target values. For the 5 m and 2.5 m models, the point dispersion becomes increasingly noticeable. Interestingly, although the 2.5 m model exhibits the largest overall spread around the identity line, it produces fewer extreme errors than the 10 m and 5 m models. In contrast, the frequency of extreme errors increases as resolution decreases.
@@
Figure 5 shows the error distribution across tree-height bins (1.5–5, 5–10, 10–15, 15–20, 20–25, 25–30, and >>30 m) using box-and-whisker plots. Overall, all three resolutions exhibit similar behavior, with a slight degradation in performance for the smallest and tallest trees. In contrast, for mid-range heights (10-15 and 15-20 m), the models perform nearly identically, with median errors close to zero. Moreover, all models tend to overestimate low height values, while for higher values they exhibit the opposite behavior. The most noticeable performance differences between the models occur above 25 m, where their accuracy drops more sharply than in the mid-range bins. This may be explained by the uneven distribution of tree heights in the dataset or by the limitations of Sentinel-2 in distinguishing fine spectral differences in short and tall forest stands -— an issue that could be addressed in future work.
[…]
@@ 4.2.2 Qualitative Results
Figure 6: $640\times 640$ m Tree height prediction patches centred on 44.893°N and 0.997°W. Subfigures (b)–(d) present the prediction results of our model at 10 m, 5 m, and 2.5 m resolutions, respectively. Subfigure (a) shows a Sentinel-2 image from the input time series acquired on 7 July 2023, while (f) depicts the reference LiDAR-derived height map at 2.5 m resolution. Subfigures (g) and (h) display the 10 and 5 m predictions upsampled to 5 m and 2.5 m respectively using bicubic interpolation. Finally, (e) provides a VHR Bing image for visual comparison. Zoomed-in views (top-left corner) highlight differences in fine details between predictions. No masking is applied in these examples.
@@
Figure 7: Normalized log-Frequency Attenuation Profile FAP for HR reference patch, SR results and bicubic interpolation for 5 m (a) and 2.5 m (b) predictions for patch presented in Figure 6. Bicubic interpolation profiles correspond to Subfigures 6(g) and 6(h) respectively. To maintain consistency, we used HR reference data with the same native resolution for both predictions. Spatial frequencies in x axis correspond to normalized spatial frequencies $f=\frac{f_{n}}{f_{N}}$, where $f_{N}$ is the maximum spatial frequency that can be represented at a given image resolution.

### v2

@@ 3.1 Model
In this work, we propose an end-to-end model, named THREASURE-Net, that performs super-resolution and regression tasks simultaneously. Additionally, our model produces a tree/non-tree classification mask, based on tree height regression features. Given a Sentinel-2 SITS $s=[s_{t_{1}},s_{t_{2}},\dots,s_{T}]\in\mathbb{R}^{T\times C\times H\times W}$ acquired at dates $t=[t_{1},t_{2},\dots,t_{T}]\in\mathbb{R}^{T}$, the model aims to predict the tree height $95^{\text{th}}$ percentile corresponding to the reference date $t_{L}$, producing an output of shape $f\cdot H\times f\cdot W$, where $f$ denotes the SR factor. Here, $T$, $C$, $H$, and $W$ represent the number of acquisition dates, spectral channels, image height, and image width in pixels, respectively. Note that for the sake of simplicity, we omit batch dimension in the model description.
@@
The RDB-FE module is extracted from Residual Dense network (RDN) (Zhang et al., 2018), which is mainly used for natural image SISR task and originally consists of four parts: shallow feature extraction net (SFENet), residual dense blocks (RDBs), dense feature fusion (DFF) and finally the up-sampling net (UPNet). Given the multi-temporal nature of our model, at this stage, we only deploy the feature extraction modules (SFENet, RDBs, and DFF). The up-sampling module is applied later to the temporally encoded features, rather than to each image individually.
@@
Temporal feature extraction block. To extract a unique set of temporal features for each Sentinel-2 series, we exploit the Lightweight Temporal Self-Attention mechanism proposed by Garnot and Landrieu (2020). The module(here referred to as Temporal Encoder) captures dependencies across the multi-temporal Sentinel-2 time series and allows the network to dynamically weight the contribution of each timestamp, by combining keys $K$, queries $Q$ and values $V$ to produce a self-attention mask.. The simplified version of this algorithm can be summarized as follows:
Let $\mathbf{f}=[\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{T}]\in\mathbb{R}^{T\times F_{1}}$ be the sequence of encoded Sentinel-2 observations at times $\{t_{1},\dots,t_{T}\}$. Each timestamp is associated with a temporal positional encoding $\mathbf{p}_{i}\in\mathbb{R}^{d}$ using a sinusoidal function from the Positional Encoding (PE) module:
@@
with $\tau$ a scaling constant and $k$ indexing the embedding dimensions. The positional encoding is added to the input features:
$\tilde{\mathbf{f}}_{i}=\mathbf{f}_{i}+\mathbf{p}_{i}.$ (2)
The attention mechanism computes keys $\mathbf{K}$ and values $\mathbf{V}$ as linear projections of $\tilde{\mathbf{f}}$, while $\mathbf{Q}$ is a learnable parameter. The resulting aggregated temporal representation has a final size $T\times F_{2}\times H\times W$, where $F_{2}$ is the feature dimension after encoding and $H\times W$ are the spatial dimensions. For full details on the Lightweight Temporal Transformer, we refer the reader to Garnot and Landrieu (2020).
@@
where $\mathbf{e}_{t_{L}}$ is encoded LiDAR dates, and $W_{Q}\in\mathbb{R}^{d\times d_{a}}$ are learnable weights.
@@
As shown in Equations 4 and 5, the resulting DOYs represent relative time offsets (in days) from the beginning and the middle of the year, respectively. This preprocessing highlights the seasonal relationship between the LiDAR reference date and the Sentinel-2 time series, allowing the model to better capture phenological phase alignment. During inference, predictions are by default conditioned on July 1st, whereas during training the DOY representation preserves temporal offsets around this date, enabling the model to encode whether an observation occurs earlier or later in the seasonal cycle.
@@
Tree height regression and classification. The final tree-height predictions and tree masks are generated by a Multi-Layer Perceptron (MLP) block (referred to as MLP regression). The MLP takes the super-resolved features as input and processes them through a set of shared layers. The final layer is task-specific, with two parallel heads. Each head performs a distinct task: the regression head predicts tree height, while the classification head outputs a tree/non-tree map at the chosen resolution. The classification head is optimized independently using detached features, preventing its gradients from affecting the shared representation.
@@ 3.2 Loss functions
To optimize the model, we employ the following loss functions: the L1 loss (Mean Absolute Error, MAE) for tree height prediction, Binary Cross-Entropy (BCE) loss for the tree/non-tree classification mask, and the Weighted Gradient Difference Loss (WGDL) to enhance the texture and edges of the super-resolved features. The total loss is defined as:
$\mathcal{L}_{total}=\omega_{height}\cdot\mathcal{L}_{height}+\omega_{classif}\cdot\mathcal{L}_{classif}+\omega_{wgdl}\cdot\mathcal{L}_{wgdl}.$ (6)
where $\omega_{height}$, $\omega_{classif}$, and $\omega_{wgdl}$ are weighting coefficients controlling the relative contribution of each loss term.
Tree height regression Loss. By design choice, the tree classification mask extraction is based on the height regression features. For this reason, during training, the height prediction loss is computed both for tree and non-tree pixels. However, the height of non-tree pixels is set to zero. Therefore, we apply weights to balance the contribution of tree and non-tree pixels, preventing the loss from being dominated by non-trees.
In this work, we assume that most patches exhibit low variability in pixel values due to the homogeneous forest structure, which may lead to imbalance within a batch. To solve this issue, instead of directly averaging pixel-wise loss values across the batch, we first average them within each patch and then across patches in the batch while applying tree/non-tree weights $w_{\text{tree}}$ and $w_{\text{non-tree}}$ loss as follows:
$\begin{split}\mathcal{L}_{\text{batch}}&=\frac{\sum_{i=1}^{N_{b}}w_{i}\,\ell(p_{b,i})}{\sum_{i=1}^{N_{b}}w_{i}},\\ \text{where }w_{i}&=\begin{cases}w_{\text{tree}}=1,&\text{if pixel }i\text{ is tree},\\ w_{\text{non-tree}},&\text{otherwise}.\end{cases}\end{split}$ (7)
$B$ is the batch size, $N_{b}$ is the number of valid pixels within a patch and $\ell(p_{b,i})$ are pixel-wise loss values. For simplicity, we set $w_{\text{tree}}$=1 for all the experiments.
@@
Moreover, for WGDL, we apply $w_{\text{non-tree}}$ weight to non-tree pixels following the same principle as for L1 loss.
@@ 4.1.1 Evaluation Metrics
To evaluate the performance of the model for tree height prediction, we report four complementary metrics: the coefficient of determination (R²), the MAE, the root mean squared error (RMSE) and the relative mean absolute error (rMAE). We additionally report Intersection over Union (IoU) to asses the quality of the tree classification mask.
R² measures how well the predicted values follow the variability of the true heights, with higher values indicating better explanatory power. MAE reflects the average magnitude of the prediction errors, while RMSE emphasizes larger errors more strongly and, therefore, highlights occasional large deviations. The rMAE expresses the average error relative to the true values, enabling comparison across height ranges with different scales. Finally, IoU quantifies the spatial overlap between predicted and reference tree masks, measuring the ratio between their intersection and union, and thus evaluates how accurately the model delineates tree-covered areas.
@@ 4.1.2 Model Configuration
We evaluate THREASURE-NET at three spatial resolutions: 10 m, 5 m, and 2.5 m. For training we use both tree and non-tree pixels, considering the pixels with values superior to 1.5 m as trees (permanent vegetation). However, the tree height regression validation metrics are exclusively computed over tree pixels derived from reference data, whereas the IoU metric is evaluated separately to assess the accuracy of tree/non-tree classification.
@@
Finally, the prediction head consists of an MLP with layer sizes (64, 128, 64) followed by two task-specific output heads: one producing a single-channel tree-height regression map and the other generating a binary classification mask.
The maximum time-series length is set to 12, and all series shorter than 5 observations are discarded. During training, a random sampling strategy is applied when the time-series length exceeds 12, whereas an equal-range sampling strategy is used during validation to ensure the consistency of the validation data.
The input Sentinel-2 patches contain $64\times 64$ pixels ($640\times 640$ m²), augmented with border margins that are passed through the model but discarded from the output before loss computation and inference, in order to avoid border effects due to 2D convolution operations. During training, we apply a random sampling strategy within the borders of each LiDAR HD patch ($1\times 1$ km²), whereas during validation, the central region of the patch is systematically selected. For this reason, the initial Sentinel-2 patches were constructed with 250 m margins on each side relative to the LiDAR patches.
As mentioned in Section 2.3, patch-weighted L1 loss, BCE and WGDL losses are used for model optimization. The loss weights were set as follow: $\omega_{height}=1$, $\omega_{classif}=1$, and $\omega_{wgdl}=0.33$. The weight for non-tree pixels is set to $w_{\text{non-tree}}=0.1$ for both patch-weighted L1 and wGDL losses.
All input features were standardized to ensure stable training.
@@
4.2 Results and Discussion
4.2.1 Quantitative Results
Table 2: Quantitative results for our models of 10, 5 and 2.5 m resolutions at the test set. We also report the performance of two other recent state-of-the-art algorithms developed for french temperate forests: FORMS-T (Schwartz et al., 2025) at 10m and Open-Canopy (Fogel et al., 2025) at 1.5m resolution. N/R stands for “Not Reported”. Here we report the performance of concurrent methods for reference purposes only.
Model Res., m MAE, m MAE, % $R^{2}$ RMSE, m IoU, %
THREASURE-Net 10 2.63 22.3 0.75 3.55 88.5
5 2.70 25.3 0.72 3.66 86.7
2.5 2.89 29.7 0.67 3.92 84.7
Open-Canopy 1.5 2.52 22.9 N/R 4.02 90.5
FORMS-T 10 3.06 N/R 0.68 N/R 76.8
*The authors report tree classification within a forest mask, while we perform tree/non-tree classification for the whole test dataset.
[…]
@@ 4.2.1 Quantitative Results
Figure 5: Prediction–target scatterplots for our 10 m, 5 m, and 2.5 m models for tree pixels. For visualization purposes, we randomly sampled an equal number of points ($\approx$96 million) from each model, corresponding to the total number of points in the 10 m dataset. The identity line is shown in white. Note that although the minimum target values in the dataset are set to 1.5 m, we report the model predictions without applying any clipping.
In this subsection, we report the quantitative results of our model performance at different resolutions for the test set. The different metrics are reported in Table 2. For reference, we report the performance of two recent state-of-the-art algorithms developed for French temperate forests, FORMS-T (Schwartz et al., 2025) and Open-Canopy (Fogel et al., 2025), using results reported in their original publications. A direct comparison with our method on independent LiDAR data is provided in Section 4.4.
@@
Our model achieves mean absolute errors of 2.63 m, 2.70 m, and 2.89 m at 2.5 m, 5 m, and 10 m resolution, respectively. Although Open-Canopy obtains the lowest MAE overall, its RMSE is higher than the RMSE achieved by THREASURE-Net at each of the three resolutions. This indicates that Open-Canopy performs well on average values but struggles more with the extreme ones, tending to regress predictions toward the mean. In our case, quantitative performance decreases as the target resolution becomes finer. Nevertheless, the 2.5 m version of our model remains competitive with Open-Canopy, despite relying exclusively on freely available 10 m Sentinel-2 imagery. By contrast, Open-Canopy is trained on 1.5 m SPOT-6/7 data, which are commercial products or accessible only through specific research agreements. This highlights an important advantage of our approach: comparable performance can be achieved without relying on costly high-resolution remote sensing input data. Moreover, all variants of THREASURE-Net reached better results than 10 m resolution FORMS-T, despite relying solely on Sentinel-2 data, whereas FORMS-T uses Sentinel-1 and Sentinel-2 data, which requires considerably heavier data processing. Figure 5 presents the prediction–target scatterplots for our models. The 10 m model produces predictions that cluster closely around the identity line, indicating strong agreement with the target values. For the 5 m and 2.5 m models, the point dispersion becomes increasingly noticeable. Interestingly, although the 2.5 m model exhibits the largest overall spread around the identity line, it produces fewer extreme errors than the 10 m and 5 m models. In contrast, the frequency of extreme errors increases as resolution decreases.
The superior performance of Open-Canopy in forest mask classification at all resolutions suggests that the classification component of our model could be further optimized.
@@
Figure 4 shows the error distribution across tree-height bins (1.5–5, 5–10, 10–15, 15–20, 20–25, 25–30, and >>30 m) using box-and-whisker plots. Overall, all three resolutions exhibit similar behavior, with a slight degradation in performance for the smallest and tallest trees. In contrast, for mid-range heights (10-15 and 15-20 m), the models perform nearly identically, with median errors close to zero. Moreover, all models tend to overestimate low height values, while for higher values they exhibit the opposite behavior. The most noticeable performance differences between the models occur above 25 m, where their accuracy drops more sharply than in the mid-range bins. This may be explained by the uneven distribution of tree heights in the dataset or by the limitations of Sentinel-2 in distinguishing fine spectral differences in short and tall forest stands – an issue that could be addressed in future work.
@@ 4.2.2 Qualitative Results
Figure 6: $640\times 640$ m tree height prediction patches centred on 44.893°N and 0.997°W. Subfigure (a) shows a Sentinel-2 image from the input time series acquired on 7 July 2023, and (b) provides a VHR Bing image for visual comparison. Subfigures (c)–(e) present the prediction results of our model at 10 m, 5 m, and 2.5 m resolutions, respectively. Subfigures (f) and (g) display the 10 m and 5 m predictions upsampled to 5 m and 2.5 m, respectively, using bicubic interpolation. Finally, (h) depicts the reference LiDAR-derived height map at 2.5 m resolution. Zoomed-in views (top-left corner) highlight differences in fine details between predictions. No masking is applied in these examples.
@@
Figure 7: Normalized log-Frequency Attenuation Profile FAP for HR reference patch, SR results obtained with THREASURE-Net and bicubic interpolation for 5 m (a) and 2.5 m (b) predictions for patch presented in Figure 6. Bicubic interpolation profiles correspond to Subfigures 6(g) and 6(h) respectively. To maintain consistency, we used HR reference data with the same native resolution for both predictions. Spatial frequencies in x axis correspond to normalized spatial frequencies $f=\frac{f_{n}}{f_{N}}$, where $f_{N}$ is the maximum spatial frequency that can be represented at a given image resolution.